T = H + 10
T + H = 42
so T = 26, H = 16
Therefore Prob(T) = 26/42 = 0.6190
If a die is rolled 3 times what is probability of having a 5 appear at least 3 times in a row?
The chance of any 1 number appearing in a die roll 3 times in a row is:
1/6 * 1/6 * 1/6
which is 1/216 = .00463
What is the mean of frequency distribution?
It is a mathematically calculated summary statistic. With discrete distributions it is the arithmetic mean whereas with a continuous distribution it is the value of the random variable (RV) such that it divides the area under the probability distribution curve in half.
1/4
Why is 'heads' plural in the phrase 'heads or tails'?
because there are many people who do this and it just sounds better. would u say head or tail? that just doesnt sound as good
Prob(Coin shows H) = 1/2
Prob(Die shows 2) = 1/6
Prob(Both events) = 1/2 * 1/6 (because they are independent)
= 1/12
Flash Back
A pair of dice is thrown Find the probability of getting a total of either 4 or 6?
There are 36 possible outcomes. Of these the following 8 meet the requirements:
(1,3), (2,2), (3,1), (1,5), (2,4), (3,3), (4,2) and (5,1).
So the probability of 4 or 6 is 8/36 or 2/9
What is an equation for a set of number 5 10 20 40 60 80?
That's easy! It's just doubling. the equation would be (insert variable here) X2
The answer would be 25%.
What is the probability that you will roll doubles on a pair of dice?
The probability that you will roll doubles on a pair of dice is 1 in 6.
The probability that you roll "something" on the first die is 1 in 1. The probability that the second die will match the first die is 1 in 6. The resultant probability is simply the product of (1 in 1) and (1 in 6).
What is the probability of getting struck by a stray bullet?
Well as someone who was hit but a stray bullet in Brooklyn, NY I can tell you depending on where you live/go the probability goes up. Im from LI, Grew up in jamica queens but was with friends in Brooklyn. In case you were saying to your self "well he must be into something illegal/wrong to get hit) NO! im not a gang/drug dealing/violence related person, in fact i am an orthodox Jewish person but was in the wrong place at wrong time (Long story, 2 unrelated people were in a fight down the block and next thing you know they started shooting at each other, and I got hit by a stray) so don't think it cant happen to you, because it can happen to anyone). BULLETS DON'T DISCRIMINATE! out there (Parts of Brooklyn) there are 8 shootings (unrelated) per week, AVERAGE (And im taking within a 3 mile radius in bed-sty) In other areas it is same, less or unfortunately more. I signed up for a daily e-mail alert from a police-related crime statistic wesite (after the shooting to get an idea of how common it was in the area i was in) and let me tell you there is a shooting every day, not to mention the other violent crimes, and if there wasn't a shooting on, say Monday, then there were 2 on Tuesday. It happened no longer than 4 months ago but it still feels like yesterday.
What if you have a wheat penny that looks like it is made out of copper but it probably isn't?
All wheat cents except those minted in 1943 were made of bronze which is 95% copper. In 1943 cents were made of steel because copper was needed for ammunition.
Please post a new question with more details about your coin, including the date and whether it has a mint mark.
Probability of not primes from 1 to 50?
A prime number is a number in N that has exactly two different numbers in N that divide it: 1 and itself.
2, 3, 5, 7, 11, 13, 17, 19, 23, 29, 31, 37, 41, 43, 47 are the prime numbers from 1 to 50
Probability of prime from 1 to 50 is 15/50 = 3/10
Probability of not prime is the complement = 1- 3/10 = 7/10
That is known as a simple random sample, or SRS.
How many different 7 letter arrangements can be formed from the letters in the word success?
There are 7!/(3!*2!) = 420 ways.
7! for the seven letters in "success", but
I dont think there is such a thing as paternal twins...perhaps you are thinking of fraternal twins. In that case, they are not identical twins, they do not look the same and they can be different sexes.
Different countries have different systems of coinage so without knowing which country the question is about, it is not possible to answer the question.
There are 4 aces in the deck the odds that the first card is an ace is 4/52 or 1/13. The odds the second card is an ace is 3/51 or 1/17 because there are only 3 aces and 51 cards left. The odds that both are aces are 1/13 times 1/17 which is 1/221.
What is the total number of outcomes for this event if Alan tosses a six sided number cube?
The total number of outcomes if Alan tosses the cube is 1 since he only tosses the cube once. However, Alan could roll a 1, 2, 3, 4, 5, or 6.
How many times must you throw two standard dice to be sure you get the same sum at least twice?
The probability of gettting a particular sum on a standard set of dice depends on which sum you are seeking. For example, the sums of 2 and 12 have a probability of 1 in 36, or about 0.0278; while the sum of 7 has a probability of 6 in 36, or 1 in 6, or about 0.167.
Specifically answering the question; it is not possible to guarantee a particular outcome in a random throw, or in a series of random throws, of the dice. You can only talk about probability. Let's take the worst case of trying to throw a 2 or a 12. Even of you throw the dice 100 times, the probability is only 0.0278100, or about 2.34 x 10-156 that you will not throw the 2 or 12; so, the probability is extremely good that you will throw a 2 or 12 in 100 throws, but it is not guaranteed. That's the thing about probability.
What are some card games to play with a regular deck of cards?
War, Poker, Slap Jack, Go fish, Speed...google it.
Risk Assessment & Analysis – This is the step where we analyze all those risks identified in the previous step. We will shortlist only those important (high impact/probability) risks and move the lower priority/possibility risks to a watch list.
you toss 3 coins what is the probability that you get exactly 2 heads given that you get at least one head?